OpenStudy (anonymous):

Which equation represents a linear function that has a slope of 4/5 and a y-intercept of –6?

OpenStudy (gabebae):

y= 4/5x - 6 this is written in slope-intercept form y=mx+b m = the slope b = the y-intercept

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you do another

OpenStudy (anonymous):

can you do another

OpenStudy (gabebae):

sure

OpenStudy (anonymous):

A linear function has a slope of -7/9 and a y-intercept of 3. How does this function compare to the linear function that is represented by the equation y+11=-7/9(x-8)? A.It has the same slope and the same y-intercept. B. It has the same slope and a different y-intercept. C. It has the same y-intercept and a different slope. D. It has a different slope and a different y-intercept.

OpenStudy (gabebae):

B. It has the same slope and a different y-intercept

OpenStudy (anonymous):

thanks again. I'm new to this and just found out how to give medals

OpenStudy (gabebae):

It's okay You should fan me too if you need help on any other questions I'm close to 100 fans

OpenStudy (anonymous):

how to do that?

OpenStudy (gabebae):

Go hover over my picture or name and click on "Become a fan" which is in blue

OpenStudy (anonymous):

yeah found out already sry i like just found out

OpenStudy (gabebae):

That's okay Welcome to openstudy :D

OpenStudy (anonymous):

one more please

OpenStudy (gabebae):

sure

OpenStudy (anonymous):

The slope of the graph of the equation y=2x-2 is 2. What is the y-intercept?

OpenStudy (anonymous):

And can you explain the process please

OpenStudy (gabebae):

y=mx+b is slope-intercept form m = the slope y = the slope so -2 is the y-intercept

OpenStudy (anonymous):

oh okay got it. I needed help cause I'm horrible in math and i made a 20% on the first attempt of the quiz and i keep messing up in the process of solving the math work so thanks again a lot.

OpenStudy (anonymous):

Now the second attempt wish me luck!

OpenStudy (gabebae):

good luck

OpenStudy (anonymous):

: / )

OpenStudy (gabebae):

how did you do?

OpenStudy (anonymous):

still on

OpenStudy (gabebae):

ok

OpenStudy (anonymous):

help

OpenStudy (gabebae):

yes?

OpenStudy (anonymous):

What are the slope and Y-intercept form of the linear function that is represented by the table. X > -3, 0, 3, 6 Y > 18, 12, 6, 0 A. The slope is –2, and the y-intercept is 6. B. The slope is –2, and the y-intercept is 12. C. The slope is 2, and the y-intercept is 6. D. The slope is 2, and the y-intercept is 12.

OpenStudy (anonymous):

this (>) is an arrow just letting you know

OpenStudy (gabebae):

Graph the points and take any 2 points and find the slope using the slope formula \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (anonymous):

still confused

OpenStudy (anonymous):

do i do X > 3 - 6 then the Y > 6 - 0

OpenStudy (anonymous):

the other way around

OpenStudy (gabebae):

no it doesn't matter?

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Is it B. The slope is –2, and the y-intercept is 12.

OpenStudy (anonymous):

oh it was right that's what my answer says. And how do you get out of this and submit another question when i want to do another

OpenStudy (anonymous):

i don't want to continue or else i'd be asking my entire quiz

OpenStudy (anonymous):

Thanks that's all